i did nt underst00d the f0II0wing Iine given in the n0tes


Let P(x1, y1) and Q(x2, y2) be two neighbouring points on the parabola y2 – 4ax. Then the equation of the line joining P and Q is

       y – y1 = (y2-y1) / (x2-x1 ) (x – x1) …… (1)


-kiran

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear kiran

we know that slope of the line joining by two point is 

 m =y2 -y1 /x2 -x1

 

now we know that equation of line passes through a point and slope   m is

 y-y1 =m( x-x1)

 now put the value of m and you will get the desired result

 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin